OpenStudy (anonymous):

what is the integral (x/(x+4) )dx ????/

myininaya (myininaya):

what if we try to let u=x+4 => du=dx if u=x+4 => x=u-4 \[\int\limits_{}^{}\frac{u-4}{u} du=\int\limits_{}^{}(1-\frac{4}{u}) du=u-4\ln|u|+C=(x+4)-4\ln|x+4|+C\]

myininaya (myininaya):

check: take derivative of antiderivative (1+0)-4*1/(x+4)+0 1-4/(x+4) [(x+4)-4]/(x+4) x/(x+4) so yay! answer checks out! :)

OpenStudy (anonymous):

um no u is X to the power of 4 :(

myininaya (myininaya):

what?

myininaya (myininaya):

we have \[\int\limits_{}^{}\frac{x}{x+4} dx=(x+4)-4\ln|x+4|+C\]

myininaya (myininaya):

this is totally correct unless you made a mistake in your question

OpenStudy (anonymous):

what is the integral (x/(1+x^4))dx